# ExaGRyPE: Numerical General Relativity Solvers Based upon the Hyperbolic PDEs Solver Engine ExaHyPE

### v2

@@ Abstract
ExaGRyPE describes a suite of solvers and solver ingredients for numerical relativity that are based upon ExaHyPE 2, the second generation of our Exascale Hyperbolic PDE Engine. Numerical relativity simulations are crucial in resolving astrophysical phenomena in strong gravitational fields and are fundamental in analyzing and understanding gravitational wave emissions. The presented generation of ExaGRyPE solves the Einstein field equations in the standard CCZ4 formulation under a 3+1 foliation and focuses on black hole space-times. It employs a block-structured Cartesian grid carrying a higher-order order Finite Difference scheme with full support of adaptive mesh refinement (AMR), it facilitates massive parallelism combining message passing, domain decomposition and task parallelism, and it supports the injection of particles into the grid as static data probes or as moving tracers. We introduce the ExaGRyPE-specific building blocks within ExaHyPE 2, and discuss its software architecture and compute-n-feel:
For this, we formalize the creation of any specific astrophysical simulation with ExaGRyPE as a sequence of lowering operations, where few abstract logical tasks are successively broken down into finer and finer tasks until we obtain an abstraction level which can directly be mapped onto a C++ executable. The overall program logic is fully specified via a domain-specific Python interface, we automatically map this logic onto a more detailed set of numerical tasks, subsequently lower this representation onto technical tasks that the underlying ExaHyPE engine uses to parallelize the application, before eventually the technical tasks in turn are mapped onto small task graphs including the actual astrophysical PDE term evaluations, initial conditions, boundary conditions, and so forth. These can be injected manually by the user, or users might instruct the solver on the most abstract user interface level to use out-of-the-box ExaGRyPE implementations. We end up with a rigorous separation of concerns which shields ExaGRyPE users from technical details and hence simplifies the development of novel physical models. We present the simulations and data for the gauge wave, static single black holes and rotating binary black hole systems, demonstrating that the code base is mature and usable. However, we also uncover domain-specific numerical challenges that need further study by the community in future work.
@@ 1 Introduction
More than one hundred years ago, Einstein wrote his famous equations of General Relativity [1]. Revealing the intrinsic coupling of space and time, those equations today lay the foundation of our understanding of the space-time geometry of the Universe. The Einstein equations have a rather compact mathematical formulation manifesting in ten highly non-linear tensor equations. Due to their complexity, analytical solutions are notoriously difficult to find once the system (initial condition) lacks strong symmetries, i.e. once we tackle setups of practical relevance. We thus need numerical techniques.
@@
The field of numerical relativity has been under active development for many decades, and numerical simulations are now omnipresent when studying astrophysical phenomena involving strong gravitational fields. By recasting the Einstein field equations, numerical relativity reduces the underlying physics to a Cauchy initial value problem, which we can solve via standard numerical schemes: The coupled space-time is cut into a sequence of three-dimensional hypersurfaces, labeled by a global temporal parameter. Every hypersurface represents one “time step”, which is a snapshot or slice through the actual space-time [2]. Despite being a well-established approach, the number of mature and actively developed open-source codes available to the community is limited (e.g., Einsteintoolkit [3], GRChombo [4], SpEC [5], LEAN [6] and Dendro-GR [7]).
Figure 1: Snapshot of two black holes orbiting around each other. They eventually merge.
@@
The software packages share some similarities and have to tackle similar technological and numerical challenges. First, numerical relativity codes have to be able to simulate a vast span of spatial scales (spatial is to be read in the sense of numerical discretization of a space-time snapshot). They have to simulate over a huge computational domain, while some features—around black holes—require a very detailed spatial resolution (Figure 1). At the same time, explicit time-stepping remains the state-of-the-art in the field, even though many astrophysical phenomena of interest such as the merger of black holes require us to simulate them over a long time span while explicit time-stepping yields tiny time-step sizes. Second, the hyperbolic–parabolic nature of Einstein’s equations in combination with many nonlinear terms implies that we have to pick our numerical techniques carefully to avoid numerical instabilities and numerical inaccuracies introducing long-term pollution of the results. Higher-order methods thus have established themselves as state-of-the-art, while several further ingredients, such as appropriate boundary conditions, have to be chosen carefully to fit the experimental challenge. Finally, the complex equations lead to complicated compute kernels involving potentially thousands of lines of code. The translation of PDE terms into code is not simple. As a result of these three characteristics, competitive numerical relativity codes have to combine sophisticated software technology, e.g., good (strong) scalability plus efficiency, with advanced numeric and a clear strategy how to master the arising code complexity.
@@
The evolving system adopted in our code is a recast version of CCZ4 (Conformal and covariant formulation of the Z4) [11], which transforms the equations into a pure first-order formulation via a set of auxiliary variables. This idea is first proposed in [10], serving as the foundation for our code. Besides this original formalism where auxiliary variables are introduced into the evolving system alongside the primary variables, ExaGRyPE comes along with a second flavor: here, auxiliary variables are derived from the primary variables in a post-processing stage after every time step. This feature enables users to implement a second-order formulation similar to, e.g., GRChombo. With lower-order and higher-order implementations of both formulations at hand, we deliver an infrastructure to investigate into the numerical details and modifications of the underlying physics. We even provide an ecosystem to couple various solvers and hence make the the numerical scheme follows the characteristics of solutions.
@@ 2 Problem Statement
$R_{ab}-\frac{1}{2}g_{ab}R+\nabla_{a}Z_{b}+\nabla_{b}Z_{a}-g_{ab}\nabla^{c}Z_{c}$
@@
$ds^{2}=-\alpha^{2}dt^{2}+\gamma_{ij}\left(dx^{i}+\beta^{i}dt\right)\left(dx^{j}+\beta^{j}dt\right),$
@@
One can construct simulations with a formulation only including $\gamma_{ij}$, the spatial component $K_{ij}$ of the extrinsic curvature $K_{ab}:=-\frac{1}{2}\mathcal{L}_{\mathbf{n}}\gamma_{ab}$, $\alpha$, $\beta^{i}$ and the $Z^{a}$ field, as they already form a complete evolving system. However, this system is unstable due to a lack of strict hyperbolicity [21]. We therefore introduce a conformal factor $\phi$ and a conformal metric $\tilde{\gamma}_{ij}$ [11] such that
$\tilde{\gamma}_{ij}:=\phi^{2}\gamma_{ij},~{}\phi=[\det(\gamma_{ij})]^{-1/6},$
and further define $\tilde{A}_{ij}$, $K$, $\Theta$, $\hat{\Gamma}^{i}$ following a similar approach in the original CCZ4 formulation [11]. Further details can be found in B.
@@ 2.1 Gauge Conditions
$(\partial_{t}-\beta^{i}\partial_{i})\alpha=-\alpha^{2}g(\alpha)K,$
@@
$\partial_{t}\alpha=\beta^{k}A_{k}-\alpha^{2}g(\alpha)\left(K-K_{0}-2c\Theta\right)$
@@
$\partial_{t}\beta^{i}$ $=$ $\beta^{k}B_{k}^{i}+fb^{i},$
$\partial_{t}b^{i}-\beta^{k}\partial_{k}b^{i}$ $=$ $\partial_{t}\hat{\Gamma}^{i}-\beta^{k}\partial_{k}\hat{\Gamma}^{i}-\eta b^{i}.$
@@ 2.4 Gravitational Wave Extraction
$\psi_{4}(t,r,\theta,\phi)=\sum^{\infty}_{l=2}\sum^{l}_{m=-l}\psi^{lm}_{4}(t,r){}_{-2}Y_{lm}(\theta,\phi),$
@@
In ExaGRyPE, the spherical integration is done by the spherical t-design scheme [28], where we measure the values of the integrated function on a set of $N$ sample points on the sphere of interest. It is shown that the average of those sample values is also the average of the integrated function itself, i.e.
@@
The method is most effective when the function of interest has a degree of freedom that is close to or lower than $t$ (i.e., it can be roughly described by a polynomial up to the degree of $t$). $N$ needs to be increased if we want higher accuracy by raising $t$. In the current version of ExaGRyPE, we choose the set of sample points of $t=43$ with $N=948$ from [29].
@@ 3.2 Numerical Solvers
$\vec{Q}(t+\delta_{t})=\vec{Q}(t)+\delta tS(\vec{Q})+\sum_{i\in d}\text{Flux}^{\pm}_{i}(\vec{Q})\Big{|}_{\partial v},$
@@
$\text{Flux}_{i}^{-}$ -=- = $\frac{\delta t}{\delta x}\Bigg{[}-\frac{1}{2}\left(\vec{Q}-\vec{Q}^{-}\right)B_{i}\left(\frac{1}{2}(\vec{Q}^{+}+\vec{Q}^{-})\right)$
@@
$\text{Flux}_{i}^{+}$ ++=+ = $\frac{\delta t}{\delta x}\Bigg{[}-\frac{1}{2}\left(\vec{Q}^{+}-\vec{Q}^{-}\right)B_{i}\left(\frac{1}{2}(\vec{Q}^{+}+\vec{Q}^{-})\right)$
@@
On this grid over sample points, we use central finite differences with a “tensor product” approach. For $K=1$, we could construct a seven-point stencil by cutting the Taylor expansion’s second-order term. In ExaGRyPE, we adopt the fourth-order accuracy scheme which requires $K\geq 2$ (we use $K=3$ in code practices for KO dissipation, see below). Its time-stepping scheme is
@@
This smoothing is often done using the so-called Kreiss-Oliger (KO) dissipation [32], which traditionally evaluates a higher-order term (a “fourth-order curvature” for example) for second-order PDEs. Its order needs to exceed that of the kernel updating scheme to avoid introducing larger numerical errors in smoothing. As we work with the fourth-order accuracy in FD4, ExaGRyPE employs the KO term at order $N=3$ in the code and makes this feed into our evaluation equation (6) with a user-defined parameter to suppress oscillations (D). The KO term can be calibrated or switched off upon demand.
@@ 3.5 Boundary Treatment
To implement a homogeneous Newmann boundary condition, we simply set $\vec{Q}_{\text{out}}=\vec{Q}_{\text{in}}$ for example. For solvers that host multiple halo layers for their time-stepping scheme, i.e. $K\geq 2$, we apply the initialization scheme to specify the values in the ghost volumes layer by layer: we move from the element layer closest to the domain boundary outwards (Figure 5).
@@
$\partial_{t}Q+c\frac{r}{x_{i}}\partial_{i}Q=-\frac{c}{r}(Q-Q_{\text{ini}})\bigg{|}_{\partial\Omega},\quad i\in\{1,2,3\}$
@@ 3.6 Adaptive Mesh Refinement
The $\vec{Q}_{\text{out}}$ data on such hanging are initialized from the coarser volumes prior to the mesh traversal. We interpolate. The counterpart is the restriction. Both schemes are implemented exclusively over face data. Alternative implementations [34] realize the interpolation over the patches themselves, i.e. use a volumetric restriction. While ExaHyPE facilitates such volumetric coupling in principle (cmp. discussion below), we do not use expensive volumetric coupling in ExaGRyPE’s use cases presently. Furthermore, we write the interpolation rules down with a tensor-product approach which ignores diagonal coupling, i.e., interpolation of elements that are not aligned along a coordinate axis, and we use exclusively a trilinear scheme.
@@
There are three open research questions to study: First, codes without 2:1 balancing are prone to strong reflections along the resolution transitions, as high-frequency waves cannot escape regions of high resolution [36]. Second, higher-order interpolation might be required to match the higher-order spatial discretion of the solvers. Finally, our resolution transition scheme does not guarantee the conservation of the interpolated and restricted physical quantities, which may unexpectedly affect the long-term evolution of the spacetime.
@@ 4.2 Programming Workflow
The project is manually lowered down into a Peano project, which is the one-to-one equivalent of C++ source code. Consequently, we dump this project then onto disk. We assume that all ExaGRyPE, Peano and ExaHyPE libraries have been built before using either the autotools or CMake. The Python script automatically picks up the used configuration arguments and creates an appropriate Makefile alongside the generated C++ code. Load balancing, additional third-party libraries, or global I/O settings all can be added to the Peano project before we write to disk.

### v3

@@ Abstract
ExaGRyPE describes a suite of solvers and solver ingredients for numerical relativity that are based upon ExaHyPE 2, the second generation of our Exascale Hyperbolic PDE Engine. Numerical relativity simulations are crucial in resolving astrophysical phenomena in strong gravitational fields and are fundamental in analyzing and understanding gravitational wave emissions. The presented generation of ExaGRyPE solves the Einstein field equations in the standard CCZ4 formulation under a 3+1 foliation and focuses on black hole space-times. It employs a block-structured Cartesian grid carrying a higher-order order Finite Difference scheme with full support of adaptive mesh refinement (AMR), it facilitates massive parallelism combining message passing, domain decomposition and task parallelism, and it supports the injection of particles into the grid as static data probes or as moving tracers. We introduce the ExaGRyPE-specific building blocks within ExaHyPE 2, and discuss its software architecture and compute-n-feel.
For this, we formalize the creation of any specific astrophysical simulation with ExaGRyPE as a sequence of lowering operations, where a few abstract logical tasks are successively broken down into finer and finer tasks until we obtain an abstraction level which can directly be mapped onto a C++ executable. The overall program logic is fully specified via a domain-specific Python interface, we automatically map this logic onto a more detailed set of numerical tasks, subsequently lower this representation onto technical tasks that the underlying ExaHyPE engine uses to parallelize the application, before eventually the technical tasks in turn are mapped onto small task graphs including the actual astrophysical PDE term evaluations, initial conditions, boundary conditions, and so forth. These can be injected manually by the user, or users might instruct the solver on the most abstract user interface level to use out-of-the-box ExaGRyPE implementations. We end up with a rigorous separation of concerns which shields ExaGRyPE users from technical details and hence simplifies the development of novel physical models. We present the simulations and data for the gauge wave, static single black holes and rotating binary black hole systems, demonstrating that the code base is mature and usable. However, we also uncover domain-specific numerical challenges that need further study by the community in future work.
@@ 1 Introduction
More than one hundred years ago, Einstein wrote his famous equations of General Relativity [1]. Revealing the intrinsic coupling of space and time, those equations today lay the foundation of our understanding of the space-time geometry of the Universe. The Einstein equations have a rather compact mathematical formulation manifesting in ten highly non-linear tensor equations. Due to their complexity, analytical solutions are notoriously difficult to find once the initial condition of the system lacks strong symmetries, i.e. once we tackle setups of practical relevance. We thus need numerical techniques.
@@
The field of numerical relativity has been under active development for many decades, and numerical simulations are now omnipresent when studying astrophysical phenomena involving strong gravitational fields. By recasting the Einstein field equations, numerical relativity reduces the underlying physics to a Cauchy initial value problem, which we can solve via standard numerical schemes: The space-time is cut into a sequence of three-dimensional hypersurfaces, labeled by a global temporal parameter. Every hypersurface represents one “time step”, which is a snapshot or slice through the actual space-time [2]. Despite being a well-established approach, the number of mature and actively developed open-source codes available to the community is limited (e.g., BAM [3], gh3d2m [4], Einsteintoolkit [5], GRChombo [6], SpEC [7], Dendro-GR [8], GR-Athena++ [9], SENR/NRPy+ [10] and SACRA [11]. A more comprehensive list can be found in Table 1 of [12]).
Figure 1: Snapshot of two black holes orbiting around each other. The colored frame represents the conformal factor field $\phi$, while the arrows show the distribution of the shift vector field $\beta^{i}$.
@@
The software packages share some similarities and have to tackle similar technological and numerical challenges. First, numerical relativity codes have to be able to simulate a vast span of spatial scales (spatial is to be read in the sense of numerical discretization of a space-time snapshot). They have to simulate over a huge computational domain, while some features—around black holes—require a very detailed spatial resolution (Figure 1). At the same time, explicit time-stepping remains the state-of-the-art in the field, even though many astrophysical phenomena of interest such as the merger of black holes require us to simulate them over a long time span while explicit time-stepping yields tiny time-step sizes. Second, numerical relativity codes usually cast Einstein’s equations into hyperbolic formulations under proper gauge choices, and its combination with many nonlinear terms implies that we have to pick our numerical techniques carefully to avoid numerical instabilities and numerical inaccuracies introducing long-term pollution of the results. Higher-order methods thus have established themselves as state-of-the-art, while several further ingredients, such as appropriate boundary conditions, have to be chosen carefully to fit the experimental challenge. Finally, the complex equations lead to complicated compute kernels involving potentially thousands of lines of code. The translation of PDE terms into code is not simple. As a result of these three characteristics, competitive numerical relativity codes have to combine sophisticated software technology, e.g., good (strong) scalability plus efficiency, with advanced numeric and a clear strategy how to master the arising code complexity.
@@
The evolving system adopted in our code is a recast version of CCZ4 (Conformal and covariant formulation of the Z4) [16], which transforms the equations into a pure first-order formulation via a set of auxiliary variables. This idea is first proposed in [15], serving as the foundation for our code. Besides this original formalism where auxiliary variables are introduced into the evolving system alongside the primary variables, ExaGRyPE comes along with a second flavor: here, auxiliary variables are derived from the primary variables in a post-processing stage after every time step. This feature enables users to implement a second-order formulation similar to many other numerical relativity codes. With lower-order and higher-order implementations of both formulations at hand, in ExaGRyPE  we implement both lower-order and higher-order schemes (solvers) for both formulations, they provide the infrastructure to investigate the numerical details and modifications of the underlying physics. We even provide an ecosystem to couple various solvers, allowing us to utilize different schemes according to the qualitative features of the solution.
@@ 2 Problem Statement
$R_{ab}-$ $\frac{1}{2}g_{ab}R+\nabla_{a}Z_{b}+\nabla_{b}Z_{a}-g_{ab}\nabla^{c}Z_{c}$ (1)
@@
One can construct simulations with a formulation only including $\gamma_{ij}$, the spatial component $K_{ij}$ of the extrinsic curvature $K_{ab}:=-\frac{1}{2}\mathcal{L}_{\mathbf{n}}\gamma_{ab}$, $\alpha$, $\beta^{i}$ and the $Z^{a}$ field, as they already form a complete evolving system. We further follow [16] and introduce a conformal factor $\phi$ and a conformal metric $\tilde{\gamma}_{ij}$, such that
$\tilde{\gamma}_{ij}:=\phi^{2}\gamma_{ij},~{}\phi=[\det(\gamma_{ij})]^{-1/6},$ (3)
and further define $\tilde{A}_{ij}$, $K$, $\Theta$, $\hat{\Gamma}^{i}$ following a similar approach in the original formulation. Further details can be found in B. We note that this formulation is just one of several candidates for a strongly hyperbolic evolution system. Other options include the generalized-harmonic formalism [26, 27], the BSSN formalism [28, 29], and other variants within the “Z4 family”, such as the original Z4 [30] and Z4c [31]. For ExaGRyPE, we have chosen to utilize CCZ4 to achieve better control over singularities and gauge conditions in black hole spacetimes.
@@ 2.1 Gauge Conditions
$(\partial_{t}-\beta^{i}\partial_{i})\alpha=-\alpha^{2}g(\alpha)K,$ (4)
@@
$\partial_{t}\alpha=\beta^{k}A_{k}-\alpha^{2}g(\alpha)\left(K-K_{0}-2c\Theta\right)$ (5)
@@
$\partial_{t}\beta^{i}$ $=$ $\beta^{k}B_{k}^{i}+fb^{i},$ (6)
$\partial_{t}b^{i}-\beta^{k}\partial_{k}b^{i}$ $=$ $\partial_{t}\hat{\Gamma}^{i}-\beta^{k}\partial_{k}\hat{\Gamma}^{i}-\eta b^{i}.$ (7)
@@ 2.4 Gravitational Wave Extraction
$\psi_{4}(t,r,\theta,\phi)=\sum^{\infty}_{l=2}\sum^{l}_{m=-l}\psi^{lm}_{4}(t,r){}_{-2}Y_{lm}(\theta,\phi),$ (13)
@@
In ExaGRyPE, the spherical integration is done by the spherical t-design scheme [38], where we interpolate the values of the integrated function from the closest grid points onto a set of $N$ sample points on the sphere of interest. It is shown that the average of those sample values is also the average of the integrated function itself, i.e.
@@
The method is most effective when the function of interest has a degree of freedom that is close to or lower than $t$ (i.e., it can be roughly described by a polynomial up to the degree of $t$). $N$ needs to be increased if we want higher accuracy by raising $t$. In the current version of ExaGRyPE, the default set of sample points is $t=43$ with $N=948$ from [39] for a balance of accuracy and performance. In practice, sets of different orders can be selected in the code, based on the accuracy requirements of individual simulations.
@@ 3.2 Numerical Solvers
$\vec{Q}(t+\delta_{t})=\vec{Q}(t)+\delta tS(\vec{Q})+\sum_{i\in d}\text{Flux}^{\pm}_{i}(\vec{Q})\Big{|}_{\partial v},$ (16)
@@
$\text{Flux}_{i}^{-}$ $=$ $\frac{\delta t}{\delta x}\Bigg{[}-\frac{1}{2}\left(\vec{Q}-\vec{Q}^{-}\right)B_{i}\left(\frac{1}{2}(\vec{Q}^{+}+\vec{Q}^{-})\right)$
@@
$\text{Flux}_{i}^{+}$ $=$ $\frac{\delta t}{\delta x}\Bigg{[}-\frac{1}{2}\left(\vec{Q}^{+}-\vec{Q}^{-}\right)B_{i}\left(\frac{1}{2}(\vec{Q}^{+}+\vec{Q}^{-})\right)$
@@
On this grid over sample points, we use central finite differences with a “tensor product” approach. For $K=1$, we could construct a seven-point stencil by cutting the Taylor expansion’s second-order term. In ExaGRyPE, we adopt the fourth-order accuracy scheme which requires $K\geq 2$ (we use $K=3$ in the practice of the code for KO dissipation, see below). Its time-stepping scheme is
@@
This smoothing is often done using the so-called Kreiss-Oliger (KO) dissipation [42], which traditionally evaluates a higher-order term (a “fourth-order curvature” for example) for second-order PDEs. Its order needs to exceed that of the kernel updating scheme to avoid introducing larger numerical errors in smoothing. As we work with the fourth-order accuracy in FD4, ExaGRyPE employs the KO term with $N=3$ (i.e., the term use value of cells up to three points away) in the code and makes this feed into our evaluation equation (19) with a user-defined parameter to suppress oscillations (D). The KO term can be calibrated or switched off upon demand.
@@ 3.5 Boundary Treatment
To implement a homogeneous Neumann boundary condition, we simply set $\vec{Q}_{\text{out}}=\vec{Q}_{\text{in}}$ for example. For solvers that host multiple halo layers for their time-stepping scheme, i.e. $K\geq 2$, we apply the initialization scheme to specify the values in the ghost volumes layer by layer: we move from the element layer closest to the domain boundary outwards (Figure 5).
@@
$\partial_{t}Q+c\frac{r}{x_{i}}\partial_{i}Q=-\frac{c}{r}(Q-Q_{\text{ini}})\bigg{|}_{\partial\Omega},\quad i\in\{1,2,3\}$ (21)
@@ 3.6 Adaptive Mesh Refinement
The $\vec{Q}_{\text{out}}$ data on such hanging faces are initialized from the coarser volumes prior to the mesh traversal. We interpolate. The counterpart is the restriction. Both schemes are implemented exclusively over face data. Alternative implementations [44] realize the interpolation over the patches themselves, i.e. use a volumetric restriction. While ExaHyPE facilitates such volumetric coupling in principle (cmp. discussion below), we do not use expensive volumetric coupling in ExaGRyPE’s use cases presently. Furthermore, we write the interpolation rules down with a tensor-product approach which ignores diagonal coupling, i.e., interpolation of elements that are not aligned along a coordinate axis, and we use exclusively a trilinear scheme.
@@
There are three open research questions to study: First, codes without 2:1 balancing are prone to strong reflections along the resolution transitions, as high-frequency waves cannot escape regions of high resolution [46]. Second, higher-order interpolation might be required to match the higher-order spatial discretion of the solvers, and we have already observed improvement in solution by using the second-order scheme in certain benchmarks below. Finally, our resolution transition scheme does not guarantee the conservation of the interpolated and restricted physical quantities, which may unexpectedly affect the long-term evolution of the spacetime.
@@ 4.2 Programming Workflow
The project is manually lowered down into a Peano project, which is the one-to-one equivalent of C++ source code. Consequently, we dump this project then onto disk. We assume that all ExaGRyPE, Peano and ExaHyPE libraries have been built before using either the autotools or CMake. The Python script automatically picks up the used configuration arguments and creates an appropriate Makefile alongside the generated C++ code. Load balancing, additional third-party libraries, or global I/O settings all can be added to the Peano project before we write to disk. While ExaHyPE allows users to define their own load-balancing strategy, the current version of ExaGRyPE utilizes the default built-in option. We plan to conduct a full investigation of load balancing strategies in future work along with the performance analysis.